OpenStudy (anonymous):

simplify 15 - 3x = 8x - 7

OpenStudy (anonymous):

solve?

OpenStudy (anonymous):

11x = = 8 x = 8/11

OpenStudy (anonymous):

15-3x=8x-7 15+7=x*(8+3) 22=11x x=2

jhonyy9 (jhonyy9):

why ??? this not is correct !!! 15-3x=8x-7 22=11x x=22/11 = 2

jhonyy9 (jhonyy9):

i have said to mr. hero !!!

OpenStudy (anonymous):

yes, my bad. johnny and olive are correct.